OpenStudy (anonymous):

(04.02 MC) A standard showerhead in Marty's house dispenses 7 gallons of water per minute. Marty changed this showerhead to an energy-saving one. The equation shows the amount of water dispensed, y, as a function of the number of minutes, x, for the new showerhead. y = 3x How much water does Marty save each day with the change in showerheads if he uses the shower for 8 minutes each day? (1 point) 4 gallons 12 gallons 32 gallons 53 gallons

OpenStudy (anonymous):

@Kit_Kat_Nat_<3 @satellite73

OpenStudy (anonymous):

instead of \(7\times 8=56\) gallons, this one uses \(3\times 8=24\)

OpenStudy (anonymous):

subtract to get your answer

OpenStudy (anonymous):

C!

OpenStudy (anonymous):

that is right, \[56-24=C\]

OpenStudy (anonymous):

thnks!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

yw